OpenStudy (anonymous):

Use a calculator to find a number "a" for which e^x>x^5 for all x>a

OpenStudy (anonymous):

\[e^x>x^5 for~all~x>a\]

OpenStudy (anonymous):

@.Sam. Do you know how to do this?

sam (.sam.):

\[e ^{x}>x ^{5}\] \[\ln(e ^{x})>\ln(x ^{5})\] \[x>5\ln(x)\] \[\frac{x}{lnx}>5\]

sam (.sam.):

Use calculator solve for x

sam (.sam.):

then the values of x is a a>0

OpenStudy (anonymous):

Ok, now how do you use the calculator in find the values of x in the first place?

sam (.sam.):

\[\frac{x}{lnx}>5\] \[\frac{x}{lnx}-5>0\] \[\frac{x-5\ln(x)}{\ln(x)}>0\] Form here, x must be >1 because if x=1, ln(1) from the denominator will be =0 which leads to infinity, x cannot be negative number because ln(-anything) is infinity, so, x>1 is the answer

OpenStudy (anonymous):

Thank you!